# Concept Conductor: Orchestrating Multiple Personalized Concepts in Text-to-Image Synthesis

### v1

@@ Introduction
Despite the excellent performance of existing methods for single-concept customization, handling multiple concepts remains challenging. Current methods(Kumari et al. 2023; Liu et al. 2023; Gu et al. 2024) often mix the attributes of multiple concepts or fail to align well with the given text prompts, especially when the target concepts are visually similar (e.g., a cat and a dog). We categorize these failures as attribute leakage and layout confusion. Layout confusion can be further divided into concept omission, subject redundancy, and appearance truncation, as shown in Figure 1. Attribute leakage denotes the application of one concept’s attributes to another (e.g., a cat acquiring the fur and eyes of a dog). Concept omission indicates one or more target concepts not appearing in the image (e.g., the absence of the target cat). Subject redundancy refers to the appearance of extra subjects similar to the target concept (e.g., an extra cat). Appearance truncation signifies the target concept’s appearance being observed only in a partial area of the subject (e.g., the upper half of a dog and the lower half of a cat).
@@
To address these challenges, we introduce Concept Conductor, a novel inference framework for multi-concept customization. This framework aims to seamlessly integrate multiple personalized concepts with accurate attributes and layout into a single image based on the given text prompts, as illustrated in Figure 2. Our method comprises three key components: multi-path sampling, layout alignment, and concept injection. Multi-path sampling allows the base model and different single-concept models to retain their independent denoising processes, thereby preventing attribute leakage between concepts. Instead of training a model containing multiple concepts to directly generate the final image, we let each single-concept model first focus on generating its corresponding concept. These generated subjects are then integrated into a single image, avoiding interference and conflict between concepts. Layout alignment ensures each model produces the correct layout, fundamentally addressing layout confusion. Specifically, we borrow the layout from a normal image and align the intermediate representations produced by each model in the self-attention layers with it. This reference image is flexible and easy to obtain. It can be a real photo, generated by advanced text-to-image models, or even a simple collage created by the user. Concept injection enables each concept to fully inject its visual features into the final generated image, ensuring harmony. We use shape-aware masks to define the generation area for each concept and inject the visual details (including structure and appearance) of personalized concepts through feature fusion in the attention layers. At each step of the denoising process, we first use layout alignment to correct the input latent space representation, and then use concept injection to obtain the next representation. Multi-path sampling is implemented in both layout alignment and concept injection to ensure the independence of each subject and coordination between different subjects.
@@
  • We propose Concept Conductor, an inference framework for multi-concept customization that utilizes multiple single-concept models to combine personalized concepts into a harmonious image, preserving visual details of each concept.
  • We develop a self-attention-based spatial guidance method to align the layout of the generated image with an easily obtainable reference image, ensuring correct layout. It maintains the diversity of the target subjects’ poses and avoids the degradation of image quality caused by coarse visual conditions.
  • Extensive experiments validate our method’s effectiveness and superiority over existing methods.
@@
Several works(Ruiz et al. 2023; Gal et al. 2022; Voynov et al. 2023; Alaluf et al. 2023) have customized text-to-image models to generate images of target concepts in new contexts by learning new visual concepts from user-provided example images. For instance, DreamBooth(Ruiz et al. 2023) embeds specific visual concepts into a pre-trained model by fine-tuning its weights. Textual Inversion(Gal et al. 2022) represents new concepts by optimizing a text embedding, later improved by P+(Voynov et al. 2023) and NeTI(Alaluf et al. 2023). Custom Diffusion(Kumari et al. 2023) explores multi-concept customization through joint training, but it requires training a separate model for each combination and often faces severe attribute leakage. Recent works(Liu et al. 2023; Gu et al. 2024) propose frameworks that combine multiple single-concept models and introduce manually defined layouts in attention maps to aid generation. For example, Cones 2(Liu et al. 2023) proposes residual embedding-based concept representations for textual composition and emphasizes or de-emphasizes a concept at a specific location by editing cross-attention. Mix-of-show(Gu et al. 2024) merges multiple custom models into one using gradient fusion and restricts each concept’s appearance area through region-controlled sampling. These works cannot fully avoid interference between concepts, leading to low success rates when handling similar concepts or more than two concepts. Additionally, by only manipulating cross-attention and neglecting the impact of self-attention on image structure, these methods often result in mismatched structures and appearances, causing layout control failures. In contrast, our method prevents attribute leakage by isolating the sampling processes of different single-concept models and achieves stable layout control through self-attention-based spatial guidance.
@@
Our method comprises three components: multi-path sampling, layout alignment, and concept injection, as illustrated in Figure 2. At each denoising step, we first correct the input latents $z_{t}$ through layout alignment, obtaining new latents $z_{t}\prime$ that carry the layout information from the reference image. Then, we inject the personalized concepts from the custom models into the base model and denoise $z_{t}\prime$ to obtain the next latents $z_{t-1}$. Multi-path sampling is implemented in both layout alignment and concept injection to ensure the independence of each concept and coordination between different subjects.
@@
Existing multi-concept customization methods often suffer from layout confusion (as shown in Figure 1), especially when the target concepts are visually similar or numerous. To address these challenges, we introduce a reference image to correct the layout during the generation process. For example, to generate an image of a specific dog and cat in a specific context, we only need a reference image containing an ordinary cat and dog. One simple approach to achieve layout control is to convert the reference image into abstract visual conditions (e.g., keypoints or sketches) and then use ControlNet(Zhang, Rao, and Agrawala 2023) or T2I-Adapter(Mou et al. 2024) for spatial guidance, which limits variability and flexibility and may reduce the fidelity of the target concepts. Another approach is to directly inject the full self-attention of the reference image into the generation process, transferring the overall structure of the image(Hertz et al. 2022). This strictly limits the poses of the target subjects, reducing the diversity of the generated images. Additionally, it requires structural similarity between the reference image subjects and target concepts to avoid distortions caused by shape mismatches.
To align the layout while preserving the structure of the target concepts, we propose a gradient-guided approach, as shown in Figure Figure 4. Given a reference image, we perform DDIM inversion to obtain the latent space representation $z_{t}^{\text{ref}}$ and record the self-attention features $F_{t}^{\text{ref}}$ at each timestep. Similarly, we record the self-attention features of the base model and each custom model during the generation process, denoted as $F_{t}^{\text{base}}$ and $F_{t}^{V_{i}}$, respectively, as shown in Figure 2. An optimization objective is set to encourage the generated image’s layout to align with the given layout:
@@
$\epsilon_{t}^{\text{fuse}}=\epsilon_{t}^{\text{base}}\odot M_{\text{base}}+\sum\limits_{i=1}^{N}\epsilon_{t}^{V_{i}}\odot M_{V_{i}}$ (3)
@@
Since the poses of the generated subjects are uncertain, predefined masks may not precisely match the shapes and contours of the target subjects, leading to incomplete appearances. To address this, we use mask refinement to allow the masks to adjust according to the shapes and poses of the target subjects during the generation process. Inspired by local-prompt-mixing(Patashnik et al. 2023), we use self-attention-based semantic segmentation to obtain the masks of the target subjects. For each target concept $V_{i}$, we cluster the self-attention of the custom model $\epsilon_{V_{i}}$ to obtain a semantic segmentation map $S_{t}^{V_{i}}$ and extract the subject’s mask $M_{t}^{V_{i},\text{custom}}$. We perform the same operation on the base model $\epsilon_{\text{base}}$, obtaining the semantic segmentation map $S_{t}^{\text{base}}$ and several masks $M_{t}^{V_{i},\text{base}}$, $i\in[1,N]$, each corresponding to a subject $V_{i}$. To reconcile the shape differences between the subjects in the base model and the custom models, the corresponding masks are merged: $M_{t}^{V_{i}}=M_{t}^{V_{i},\text{custom}}\cup M_{t}^{V_{i},\text{base}}$.
For initialization, we perform DDIM inversion on the reference image and extract the original masks $M_{T}^{V_{i}}$ from the self-attention layers in the same way. An alternative way is to use grounding models and segmentation models to extract masks in the pixel space, which provides higher resolution masks but requires additional computation and storage overhead. Through concept injection, we ensure the harmony of the image while fully preserving the attributes of the target concepts.
@@
We compare our method with three multi-concept customization methods: Custom Diffusion, Cones 2, and Mix-of-Show. For Custom Diffusion, we use the diffusers(von Platen et al. 2022) version implementation, while for the other methods, we use their official code implementations. All experimental settings followe the official recommendations. For Cones 2 and Mix-of-Show, grounding models are used to extract bounding boxes of target subjects from the layout reference image to ensure consistent spatial conditions. To ensure a fair comparison, no additional control models like ControlNet or T2I-Adapter are used.
@@
We evaluate multi-concept customization methods from two perspectives: text alignment and image alignment. For text alignment, we report results on CLIP(Radford et al. 2021) and ImageReward(Xu et al. 2024). For image alignment, we introduce a new metric called “SegSim” to address the limitations of traditional image similarity methods, which cannot reflect attribute leakage and layout conflicts. SegSim evaluates fine-grained fidelity by using text-guided grounding models and segmentation models to extract subject segments from generated and reference images, then calculating their similarity. Detailed information is in Appendix A.3. We use CLIP(Radford et al. 2021) and DINOv2(Oquab et al. 2023) to calculate segment similarity and report image alignment based on these models. To systematically evaluate omission and redundancy in multi-concept generation, grounding models are used to automatically count the number of target category subjects in each generated image.
@@
To verify the effectiveness of the proposed components, we conduct qualitative comparisons of various settings, as shown in Figure 8. In Figure 8(a), removing layout alignment lead to incorrect layouts, including appearance truncation (e.g., two dogs incorrectly stitched together) and concept omission (e.g., missing turquoise cup). Figures 8(b) and 8(c) show that disabling either self-attention or cross-attention features during concept injection results in a decline in fidelity, indicating that both are crucial for preserving the visual details of the target concepts. Figure 8(d) demonstrates that ablating mask refinement can cause generated subjects’ contours to not match the target concepts, leading to incomplete appearances (e.g., chow chow’s fur, corgi’s ears, green cup’s handle). To avoid randomness, we conduct quantitative comparisons of various settings using the same data and evaluation metrics as in the previous section, with results reported in Table 2. As shown in Table 2, layout alignment effectively avoids omission and redundancy, significantly improving the alignment of generated images with textual semantics. Feature fusion in both self-attention and cross-attention layers leads to higher image alignment, as both are crucial for reproducing the attributes of the target concepts. Mask refinement further improves text alignment and image alignment by optimizing the edge details of the generated subjects.

### v3

@@ Introduction
Despite the excellent performance of existing methods for single-concept customization, handling multiple concepts remains challenging. Current methods(Kumari et al. 2023; Liu et al. 2023b; Gu et al. 2024) often mix the attributes of multiple concepts or fail to align well with the given text prompts, especially when the target concepts are visually similar (e.g., a cat and a dog). We categorize these failures as attribute leakage and layout confusion. Layout confusion can be further divided into concept omission, subject redundancy, and appearance truncation, as shown in Figure 1. Attribute leakage denotes the application of one concept’s attributes to another (e.g., a cat acquiring the fur and eyes of a dog). Concept omission indicates one or more target concepts not appearing in the image (e.g., the absence of the target cat). Subject redundancy refers to the appearance of extra subjects similar to the target concept (e.g., an extra cat). Appearance truncation signifies the target concept’s appearance being observed only in a partial area of the subject (e.g., the upper half of a dog and the lower half of a cat).
@@
To address these challenges, we introduce Concept Conductor, a novel inference framework for multi-concept customization. This framework aims to seamlessly integrate multiple personalized concepts with accurate attributes and layout into a single image based on the given text prompts, as illustrated in Figure 2. Our method comprises three key components: multipath sampling, layout alignment, and concept injection. Multipath sampling allows the base model and different single-concept models to retain their independent denoising processes, thereby preventing attribute leakage between concepts. Instead of training a model containing multiple concepts to directly generate the final image, we let each single-concept model first focus on generating its corresponding concept. These generated subjects are then integrated into a single image, avoiding interference and conflict between concepts. Layout alignment ensures each model produces the correct layout, fundamentally addressing layout confusion. Specifically, we borrow the layout from a normal image and align the intermediate representations produced by each model in the self-attention layers with it. This reference image is flexible and easy to obtain. It can be a real photo, generated by advanced text-to-image models, or even a simple collage created by the user. Concept injection enables each concept to fully inject its visual features into the final generated image, ensuring harmony. We use shape-aware masks to define the generation area for each concept and inject the visual details (including structure and appearance) of personalized concepts through feature fusion in the attention layers. At each step of the denoising process, we first use layout alignment to correct the input latent space representation, and then use concept injection to obtain the next representation. Multipath sampling is implemented in both layout alignment and concept injection to ensure the independence of each subject and coordination between different subjects.
@@
  • We introduce Concept Conductor, a novel framework for multi-concept customization, preventing attribute leakage and layout confusion through multipath sampling and self-attention-based spatial guidance.
  • We develop a concept injection technique, utilizing shape-aware masks and feature fusion to ensure harmony and visual fidelity in multi-concept image generation.
  • We construct a new dataset containing 30 personalized concepts across representative categories. Comprehensive experiments on this dataset demonstrate the superior concept fidelity and alignment with textual semantics of the proposed Concept Conductor, including for concepts that are visually similar.
@@
Several works(Ruiz et al. 2023; Gal et al. 2022; Voynov et al. 2023; Alaluf et al. 2023) have customized text-to-image models to generate images of target concepts in new contexts by learning new visual concepts from user-provided example images. For instance, DreamBooth(Ruiz et al. 2023) embeds specific visual concepts into a pre-trained model by fine-tuning its weights. Textual Inversion(Gal et al. 2022) represents new concepts by optimizing a text embedding, later improved by P+(Voynov et al. 2023) and NeTI(Alaluf et al. 2023). Custom Diffusion(Kumari et al. 2023) explores multi-concept customization through joint training, but it requires training a separate model for each combination and often faces severe attribute leakage. Recent works(Liu et al. 2023b; Gu et al. 2024) propose frameworks that combine multiple single-concept models and introduce manually defined layouts in attention maps to aid generation. For example, Cones 2(Liu et al. 2023b) proposes residual embedding-based concept representations for textual composition and emphasizes or de-emphasizes a concept at a specific location by editing cross-attention. Mix-of-show(Gu et al. 2024) merges multiple custom models into one using gradient fusion and restricts each concept’s appearance area through region-controlled sampling. These works cannot fully avoid interference between concepts, leading to low success rates when handling similar concepts or more than two concepts. Additionally, by only manipulating cross-attention and neglecting the impact of self-attention on image structure, these methods often result in mismatched structures and appearances, causing layout control failures. In contrast, our method prevents attribute leakage by isolating the sampling processes of different single-concept models and achieves stable layout control through self-attention-based spatial guidance.
@@
Our method comprises three components: multipath sampling, layout alignment, and concept injection, as illustrated in Figure 2. At each denoising step, we first correct the input latents $z_{t}$ through layout alignment, obtaining new latents $z_{t}\prime$ that carry the layout information from the reference image. Then, we inject the personalized concepts from the custom models into the base model and denoise $z_{t}\prime$ to obtain the next latents $z_{t-1}$. Multipath sampling is implemented in both layout alignment and concept injection to ensure the independence of each concept and coordination between different subjects.
@@
Existing multi-concept customization methods often suffer from layout confusion (as shown in Figure 1), especially when the target concepts are visually similar or numerous. To address these challenges, we introduce a reference image to correct the layout during the generation process. For example, to generate an image of a specific dog and cat in a specific context, we only need a reference image containing an ordinary cat and dog. One simple approach to achieve layout control is to convert the reference image into abstract visual conditions (e.g., keypoints or sketches) and then use ControlNet(Zhang, Rao, and Agrawala 2023) or T2I-Adapter(Mou et al. 2024) for spatial guidance, which limits variability and flexibility and may reduce the fidelity of the target concepts. Another approach is to directly inject the full self-attention of the reference image into the generation process, transferring the overall structure of the image(Hertz et al. 2022; Kwon et al. 2024). This strictly limits the poses of the target subjects, reducing the diversity of the generated images. Additionally, it requires structural similarity between the reference image subjects and target concepts to avoid distortions caused by shape mismatches.
To align the layout while preserving the structure of the target concepts, we propose a gradient-guided approach, as shown in Figure 4. Given a reference image, we perform DDIM inversion to obtain the latent space representation $z_{t}^{\text{ref}}$ and record the self-attention features $F_{t}^{\text{ref}}$ at each timestep. Similarly, we record the self-attention features of the base model and each custom model during the generation process, denoted as $F_{t}^{\text{base}}$ and $F_{t}^{V_{i}}$, respectively, as shown in Figure 2. An optimization objective is set to encourage the generated image’s layout to align with the given layout:
@@
$\epsilon_{t}^{\text{fuse}}=\epsilon_{t}^{\text{base}}\odot M^{\text{base}}+\sum\limits_{i=1}^{N}\epsilon_{t}^{V_{i}}\odot M^{V_{i}}$ (3)
@@
Since the poses of the generated subjects are uncertain, predefined masks may not precisely match the shapes and contours of the target subjects, leading to incomplete appearances. To address this, we use mask refinement to allow the masks to adjust according to the shapes and poses of the target subjects during the generation process. Inspired by local-prompt-mixing(Patashnik et al. 2023), we use self-attention-based semantic segmentation to obtain the masks of the target subjects. For each target concept $V_{i}$, we cluster the self-attention $A_{t}^{V_{i}}$ of the custom model $\epsilon_{V_{i}}$ to obtain a semantic segmentation map $S_{t}^{V_{i}}$ and extract the subject’s mask $M_{t}^{V_{i},\text{custom}}$. We perform the same operation on the self-attention $A_{t}^{\text{base}}$ of the base model $\epsilon_{\text{base}}$, obtaining the semantic segmentation map $S_{t}^{\text{base}}$ and several masks $M_{t}^{V_{i},\text{base}}$, $i\in[1,N]$, each corresponding to a subject $V_{i}$. To reconcile the shape differences between the subjects in the base model and the custom models, the corresponding masks are merged: $M_{t}^{V_{i}}=M_{t}^{V_{i},\text{custom}}\cup M_{t}^{V_{i},\text{base}}$.
For initialization, we perform DDIM inversion on the reference image and extract the original masks $M_{T}^{V_{i}}$ from the self-attention layers in the same way. An alternative way is to use grounding models(e.g., Grounding DINO(Liu et al. 2023a)) and segmentation models(e.g., SAM(Kirillov et al. 2023)) to extract masks in the pixel space, which provides higher resolution masks but requires additional computation and storage overhead. Through concept injection, we ensure the harmony of the image while fully preserving the attributes of the target concepts.
@@
We compare our method with three multi-concept customization methods: Custom Diffusion, Cones 2, and Mix-of-Show. For Custom Diffusion, we use the diffusers(von Platen et al. 2022) version implementation, while for the other methods, we use their official code implementations. All experimental settings follow the official recommendations. For Cones 2 and Mix-of-Show, grounding models are used to extract bounding boxes of target subjects from the layout reference image to ensure consistent spatial conditions. To ensure a fair comparison, no additional control models like ControlNet or T2I-Adapter are used.
@@
We evaluate multi-concept customization methods from two perspectives: text alignment and image alignment. For text alignment, we report results on CLIP(Radford et al. 2021) and ImageReward(Xu et al. 2024). For image alignment, we introduce a new metric called Segmentation Similarity (SegSim) to address the limitations of traditional image similarity methods, which cannot reflect attribute leakage and layout conflicts. SegSim evaluates fine-grained fidelity by using text-guided grounding models and segmentation models to extract subject segments from generated and reference images, then calculating their similarity. Detailed information is in Appendix A.3. We use CLIP(Radford et al. 2021) and DINOv2(Oquab et al. 2023) to calculate segment similarity and report image alignment based on these models. To systematically evaluate omission and redundancy in multi-concept generation, grounding models are used to automatically count the number of target category subjects in each generated image.
@@
To verify the effectiveness of the proposed components, we conduct qualitative comparisons of various settings, as shown in Figure 8. In Figure 8(a), removing layout alignment leads to incorrect layouts, including appearance truncation (e.g., two dogs incorrectly stitched together) and concept omission (e.g., missing turquoise cup). Figures 8(b) and 8(c) show that disabling either self-attention or cross-attention features during concept injection results in a decline in fidelity, indicating that both are crucial for preserving the visual details of the target concepts. Figure 8(d) demonstrates that ablating mask refinement can lead to a mismatch between the generated subject contours and the target concepts, resulting in incomplete appearances (e.g., chow chow’s fur, corgi’s ears, green cup’s handle). To avoid randomness, we conduct quantitative comparisons of various settings using the same data and evaluation metrics as in the previous section, with results reported in Table 2. As shown in Table 2, layout alignment effectively avoids omission and redundancy, significantly improving the alignment of generated images with textual semantics. Feature fusion in both self-attention and cross-attention layers leads to higher image alignment, as both are crucial for reproducing the attributes of the target concepts. Mask refinement further improves text alignment and image alignment by optimizing the edge details of the generated subjects.